if {x} and [x] represent fractional and intergral part of x , then find the value of


[x]+ 2000∑r=1{x+r}/2000

Grade:

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi anurag,

[x]+ 2000∑r=1{x+r}/2000

since {1+x}={x}

{2+x}={x}

{3+x}={x}

.

.

.

.

{r+x}={x}

 the equation becomes,[x]+{x}×2000/2000 =[x] + {x} =x